# Redefining Machine Unlearning: A Conformal Prediction-Motivated Approach

### v1

@@ Abstract
Machine unlearning seeks to systematically remove specified data from a trained model, effectively achieving a state as though the data had never been encountered during training. While metrics such as Unlearning Accuracy (UA) and Membership Inference Attack (MIA) provide a baseline for assessing unlearning performance, they fall short of evaluating the completeness and reliability of forgetting. This is because the ground truth labels remain potential candidates within the scope of uncertainty quantification, leaving gaps in the evaluation of true forgetting. In this paper, we identify critical limitations in existing unlearning metrics and propose enhanced evaluation metrics inspired by conformal prediction. Our metrics can effectively capture the extent to which ground truth labels are excluded from the prediction set. Furthermore, we observe that many existing machine unlearning methods do not achieve satisfactory forgetting performance when evaluated with our new metrics. To address this, we propose an unlearning framework that integrates conformal prediction insights into Carlini & Wagner adversarial attack loss. Extensive experiments on the image classification task demonstrate that our enhanced metrics offer deeper insights into unlearning effectiveness, and that our unlearning framework significantly improves the forgetting quality of unlearning methods.
@@ 1 Introduction
Machine unlearning has become crucial in advancing data privacy, especially under legal requirements like the General Data Protection Regulation (GDPR) (Bourtoule et al., 2021). These regulations emphasize the right for individuals to have their data removed or forgotten, creating a demand for machine unlearning methods that can enable machine learning models to behave as if specific forget data were never used in training stage. The existing post hoc machine unlearning methods can be categorized into training-based (Graves et al., 2021; Warnecke et al., 2021; Thudi et al., 2022; Tarun et al., 2023) and training-free (Guo et al., 2019; Sekhari et al., 2021; Nguyen et al., 2020; Golatkar et al., 2020b, 2021; Foster et al., 2024) approaches, depending on whether they require any traditional model training steps during the unlearning process (Foster et al., 2024).
Table 1: Grad-CAM maps of one original model in CIFAR-10 with ResNet18 and two corresponding unlearning models. The Prediction row indicates whether the model correctly predicts the image’s true label, while the In Set row represents whether the true label is included in the prediction set. Although the Finetune unlearning method, can misclassify the forget data, Grad-CAM can still highlight key features of the object under this model since the true label is included in the prediction set. In contrast, our unlearning method removes the true label from the set, with activation regions shifting significantly away from the object’s key features. This confirms that the forgetting quality is better if the true label can be excluded from the prediction set.
Class Name Forget Data Original Model Finetune Method Our Method
Wok [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Swimming Trunks [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction
In Set
@@
To measure the forgetting quality and predictive performance of the model after the unlearning process (i.e., unlearning model), several unlearning metrics have been proposed (Kashef, 2021; Shokri et al., 2017; Chen et al., 2021; Brophy & Lowd, 2021; Cao & Yang, 2015). However, existing unlearning metrics, such as unlearning accuracy (UA) (Brophy & Lowd, 2021; Foster et al., 2024) and membership inference attack (MIA) (Shokri et al., 2017; Chen et al., 2021), fall short in fully evaluating forgetting quality and reliability. These metrics primarily focus on whether models can predict forget data accurately without sufficiently addressing how well the model forgets. In a nutshell, misclassifying forget data does not mean that the model has completely forgotten it to some extent.
@@
To verify this view, we apply conformal prediction (Papadopoulos et al., 2002; Lei & Wasserman, 2014) to the unlearning models. Through experiments, we find that although the model misclassifies some forget data, a significant portion of these misclassified instances still appear in the conformal prediction set. Pictures in Table 1, which visualize the important feature of models’ prediction of models by using Grad-CAM (Selvaraju et al., 2017), further explain this phenomenon. Despite the Finetune method incorrectly predicting the forget data, the Grad-CAM maps still focus heavily on the important features of the object itself.
@@
Based on the above findings, we design two novel metrics that capture the uncertainty and robustness of unlearning performance more effectively inspired by conformal prediction. Additionally, motivated by Carlini & Wagner (C&W) attack (Carlini & Wagner, 2017) and conformal prediction, we propose a general unlearning framework to improve training-based unlearning methods and promote reliable forgetting. Grad-CAM maps of our method in Table 1 reveal that once the true label no longer falls within the conformal prediction set, the activation regions shift significantly. To sum up, the contributions of our work are as follows:
  • We identify pivotal limitations in current unlearning metrics, as they overlook misclassified data where ground truth labels remain potential candidates under uncertainty quantification.
@@
  • We propose a general unlearning framework for training-based machine unlearning methods motivated by conformal prediction and C&W loss.
  • Extensive experiments demonstrate the effectiveness of novel metrics and our unlearning framework.
@@ 2.1 Preliminaries and Notations
Machine unlearning is the targeted removal of certain training data effects from a machine learning model. In our work, two different forgetting scenarios are considered: (i) random data forgetting focuses on randomly forgetting specific instances within the training data, and (ii) class-wise forgetting aims to remove all information associated with an entire class. Let $\mathcal{D}_{train}$ denote the original training data used to obtain an original model ${\bm{\theta}}_{o}$. We split the whole training data $\mathcal{D}_{train}$ into two subsets, forget data $\mathcal{D}_{f}$ and retain data $\mathcal{D}_{r}=\mathcal{D}_{train}\setminus\mathcal{D}_{f}$. In random data forgetting, $\mathcal{D}_{test}$ represents test data. In class-wise forgetting, $\mathcal{D}_{tf}$ corresponds to the test-forget data, exclusively containing the forget class, while $\mathcal{D}_{tr}$ represents the test-retain data within the test data $\mathcal{D}_{test}$. Let $\theta_{u}$ denote the model after the unlearning process, where the influence of forget data $\mathcal{D}_{f}$ has been removed. The purpose of machine unlearning is to enable $\theta_{u}$ to forget $\mathcal{D}_{f}$ (thus no longer recognizing $\mathcal{D}_{tf}$ in the class-wise forgetting scenario), while maintaining accuracy on the retain data $\mathcal{D}_{r}$ and avoiding a significant drop in accuracy on $\mathcal{D}_{test}$ in the random data forgetting scenario or $\mathcal{D}_{tr}$ in the class-wise setting.
@@
Conformal prediction is proposed to quantify uncertainty in machine learning models, providing prediction sets that contain the true label with a guaranteed probability (Angelopoulos & Bates, 2021). Among the various types of conformal prediction, this work specifically focuses on split conformal prediction (SCP) [^1] Note that while the goal is to remove the influence of the forget data so that it behaves similarly to the calibration data, the exchangeability property may not always hold in machine unlearning settings. Here, we are directly leveraging the concept of conformal prediction to evaluate machine unlearning performance.. To construct the conformal prediction set, it involves four steps:
@@
    Calibration Data. SCP first chooses unseen data as calibration data. The number of calibration data points should be enough to evaluate the model’s uncertainty.
@@
    where $p_{i}({\bm{x}})$ represents the probability output (through a softmax function) for the class $y_{i}$.
@@
    Quantile Computation. Given a target miscoverage rate $\alpha\in[0,1]$, SCP obtains threshold $\hat{q}$ by taking the $1-\alpha$ quantile of the non-conformity score on the calibration data $\mathcal{D}_{c}$.
@@
    where $y_{i}$ represents different classes.
2.2 Identifying Limitations in Current Unlearning Metrics
Evaluating the effectiveness of machine unlearning involves a trade-off between how well the model forgets unwanted data (forgetting quality) and how much it retains useful knowledge (predictive performance). The first key question we pose is as follows:
(Q1) Are existing commonly used metrics sufficient to evaluate unlearning performance?
@@ 2.2 Identifying Limitations in Current Unlearning Metrics
We apply the most commonly used metrics, unlearning accuracy (UA), retain accuracy (RA), test accuracy (TA) and MIA, to evaluate three classic machine unlearning methods: Retrain (RT), Finetune (FT) (Warnecke et al., 2021), and Random Label (RL) (Graves et al., 2021). These methods are implemented on a ResNet-18 model trained with the CIFAR-10 dataset in a random data forgetting scenario. In Table 2, compared with RA on $\mathcal{D}_{r}$, the UA of the three methods proves that the model indeed forgets some forget data. The MIA column also indicates that many data points seem to be forgotten. These metrics appear to provide valuable insights into the extent of forgetting achieved. However, higher unlearning accuracy cannot ensure that these forget data points are no longer represented in any form within the model’s predictions.
To validate our hypothesis, we employ the process of conformal prediction to examine whether the true labels of the forget data appear within the prediction set. We set the confidence level to $95\%$, i.e., $\alpha=0.05$, and set the calibration set size to $2,000$. In Table 3, we count both UA metric and MIA metric’s misclassified data points (mis-label) and, after applying conformal prediction, determined how many of these points fall within the conformal prediction set (in-set). Table 3 demonstrates that even though the model misclassifies part of the forget data, on average 54.6% of these misclassified instances are still present in the conformal prediction set. This indicates that a high UA does not mean the model has truly forgotten the data. Relying solely on UA to evaluate the forgetting quality is insufficient.
In MIA, the prediction of ‘0’ indicates a data point is predicted as “truly forget”, while ‘1’ signifies “still retain”. And “mis-label” refers to the number of the model predicts the forget data points as ‘0’. In conformal prediction of MIA, there are three potential prediction set results, i.e., $\{0\},\{1\},\{0,1\}$. The “in-set” here refers to the number of mislabeled data points whose conformal prediction set includes $1$. Thus, the in-set ratio indicates that, although the MIA fails to identify an average of 18.33% of the forget data points, conformal prediction can still recognize more than half of these forget data points within its prediction set.
Overall, high in-set ratios in conformal prediction highlight that a considerable portion of the forget data remains identifiable through conformal prediction. This implies these unlearning methods may not be fully effective in eliminating traces of the forget data from another perspective.
@@ 2.3 Designing Metrics Motivated by Conformal Prediction
Based on the limitations of UA and MIA metrics shown in Section 2.2, it raises a question as follows:
(Q2) Can we develop metrics to address the shortcomings of UA and MIA?
@@ 2.3.1 Definition of New Metrics
To overcome the limitations of UA, we introduce a novel metric, Conformal Ratio (CR), which incorporates both coverage and set size in conformal prediction to provide a more comprehensive evaluation. Before defining CR, we introduce Coverage and Set Size.
The definition of Coverage, given a dataset $\mathcal{D}$, is as follows:
@@
where $y_{t}$ is the true label of data point ${\bm{x}}$. Coverage reflects the probability that the true label falls within the prediction set ${\mathbb{C}}({\bm{x}})$. For $\mathcal{D}=\mathcal{D}_{f}$, high coverage indicates that the model retains significant information about forget data, suggesting incomplete unlearning. Conversely, higher coverage is better for test data $\mathcal{D}_{test}$.
Given a dataset $\mathcal{D}$, Set Size can be defined as follows:
@@
where $|{\mathbb{C}}({\bm{x}})|$ represents the prediction set size of a specific instance. When $\mathcal{D}=\mathcal{D}_{f}$, a smaller set size implies that the model cannot forget specific information about $\mathcal{D}_{f}$ well. In contrast, when $\mathcal{D}=\mathcal{D}_{test}$, a small set size may indicate that the model maintains a strong performance after unlearning.
@@
CR balances the information captured by coverage and set size. A lower CR value implies successful unlearning on forget data $\mathcal{D}_{f}$, but a significant drop in the model’s utility for $\mathcal{D}_{test}$. CR is inspired by conformal prediction which is proposed to assess the model’s behavior on new, unseen data (e.g., validation or test data), not on the training data. Thus, we emphasize that CR only measures forget data $\mathcal{D}_{f}$ and test data $\mathcal{D}_{test}$.
@@
To address the limitations of the current MIA metric, we propose a new metric called MIACR. For three potential prediction sets {0}, {1} and {0, 1}, set {0} is the only ideal case for MIA prediction since ‘1’ represents the data point is part of the training data. Therefore, we introduce a new metric to measure the probability of the set {0} occurring, referred to as MIACR:
@@
where ${\mathbb{C}}({\bm{x}})=\{0\}$ denotes prediction set is $\{0\}$. MIACR only calculates the average of ${\mathbb{C}}({\bm{x}})=\{0\}$, and a higher MIACR value represents better forgetting for $\mathcal{D}_{f}$.
@@
Unlike existing metrics, both the CR and MIACR metrics incorporate the fact that the true labels of some forgotten data points may remain within the prediction set, even when the predictions deviate from the true labels. This allows for a more robust assessment of the model’s forgetting quality and predictive performance.
Evaluation Criteria.
@@
Similar to existing unlearning metrics, we consider two different criteria to measure unlearning performance with our metrics. ➊ Gap to RT Criterion: A lower gap to RT method is better for both CR and MIACR metrics. The performance gap relative to the RT method is represented in (). ➋ Limit-Based Criterion: For the CR metric, a lower CR value of forget data $\mathcal{D}_{f}$ indicates better unlearning performance while a higher CR value of $\mathcal{D}_{test}$ is desirable. For the MIACR metric, a higher MIACR value for forget data ($\mathcal{D}_{f}$) reflects improved unlearning effectiveness.
2.3.2 Confidence Level and Calibration Data
Figure 1: The stability of $\hat{q}$ in different calibration set sizes (results of RT method on CIFAR-10 with ResNet-18). When the calibration set size is greater than $2,000$, the fluctuation of $\hat{q}$ can be kept within a certain range.

### v2

@@ Abstract
Machine unlearning seeks to remove the influence of specified data from a trained model. While metrics such as unlearning accuracy (UA) and membership inference attack (MIA) provide baselines for assessing unlearning performance, they fall short of evaluating the forgetting reliability. In this paper, we find that the data misclassified across UA and MIA still have their ground truth labels included in the prediction set from the uncertainty quantification perspective, which raises a fake unlearning issue. To address this issue, we propose two novel metrics inspired by conformal prediction that more reliably evaluate forgetting quality. Building on these insights, we further propose a conformal prediction-based unlearning framework that integrates conformal prediction into Carlini & Wagner adversarial attack loss, which can significantly push the ground truth label out of the conformal prediction set. Through extensive experiments on image classification task, we demonstrate both the effectiveness of our proposed metrics and the superiority of our unlearning framework, which improves the UA of existing unlearning methods by an average of $6.6\%$ through the incorporation of a tailored loss term alone.
@@ 1 Introduction
Machine unlearning has become crucial in advancing data privacy, especially under legal requirements like the General Data Protection Regulation (GDPR) Bourtoule et al. (2021). These regulations emphasize the right for individuals to have their data removed, creating a demand for unlearning methods that can enable models to behave as if specific data were never used in training stage. Beyond privacy considerations, machine unlearning has also emerged as a promising tool to remove harmful biases and stereotypes embodied in the model. The existing post hoc machine unlearning methods can be categorized into training-based Graves et al. (2021); Tarun et al. (2023); Thudi et al. (2022); Warnecke et al. (2021) and training-free Foster et al. (2024); Golatkar et al. (2021, 2020); Guo et al. (2019); Nguyen et al. (2020); Sekhari et al. (2021) approaches, depending on whether they require any model training steps during the unlearning process Foster et al. (2024).
@@
To measure the forgetting quality and predictive performance of an unlearning model, several unlearning metrics have been proposed Brophy and Lowd (2021); Cao and Yang (2015); Chen et al. (2021); Kashef (2021); Shokri et al. (2017). However, existing unlearning metrics, such as unlearning accuracy (UA) and membership inference attack (MIA), fall short in fully evaluating forgetting reliability. These metrics primarily focus on whether models can predict forget data accurately without sufficiently addressing how well the model forgets the data. In a nutshell, misclassifying forget data does not mean that the model has completely forgotten it to some extent.
To verify this view, conformal prediction Lei and Wasserman (2014); Papadopoulos et al. (2002) as an uncertainty quantification technique, is applied in our work to recover the misclassified data in UA and MIA. Through extensive experiments,
Figure 1: Grad-CAM maps of one original model and two corresponding unlearning models in CIFAR-10 with ResNet18. The Classification row indicates whether the model correctly predicts the image’s true label, while the In Set row represents whether the true label is included in the prediction set. Although the Finetune method, can misclassify the forget data, Grad-CAM can still highlight key features of the object since the true label is included in the prediction set. In contrast, when our unlearning method removes the true label from the set, activation regions shift significantly away from the object’s key features. This confirms that the forgetting quality improves if the true label can be excluded from the prediction set.
@@
we find that although the model misclassifies some forget data from the UA and MIA perspective, over $50\%$ of these misclassified data instances still appear in the conformal prediction set and can be easily recovered, which exposes a fake unlearning issue. As shown in Figure 1, the important features of models’ prediction by Grad-CAM Selvaraju et al. (2017) visualize this fake unlearning issue. Despite the Finetune method misclassifying the forget data, the Grad-CAM maps still focus heavily on the important features of the object itself.
@@
Based on the above insights, we design two novel metrics CR and MIACR that more effectively capture the uncertainty and robustness of unlearning performance inspired by conformal prediction to tackle the fake unlearning issue. Additionally, motivated by conformal prediction insights about fake unlearning and Carlini & Wagner (C&W) attack loss Carlini and Wagner (2017), we propose a general unlearning framework, which can improve existing training-based unlearning methods and promote reliable forgetting. Grad-CAM maps of our method in Figure 1 reveal that once the true label no longer falls within the conformal prediction set, the activation regions shift significantly. To sum up, our contributions are as follows:
  • Our analysis reveals that conformal prediction can recover a substantial portion of data previously classified as forgotten by existing unlearning metrics. This fake unlearning issue underscores critical limitations in existing unlearning evaluation methodologies.
@@
  • We propose an unlearning framework motivated by conformal prediction and C&W loss, enhancing existing training-based unlearning methods over both existing and our metrics.
@@ 2.1 Preliminaries and Notations
In our work, we focus on the image classification task, which is widely used in prior unlearning literature Foster et al. (2024); Graves et al. (2021); Shen et al. (2024). Two forgetting scenarios are considered in this work: (i) random data forgetting focuses on randomly forgetting specific data instances within the training data, and (ii) class-wise forgetting aims to remove all data information associated with an entire class. Let $\mathcal{D}_{train}$ denote the original training data used to obtain an original model ${\bm{\theta}}_{o}$. We split the whole training data $\mathcal{D}_{train}$ into two subsets, forget data $\mathcal{D}_{f}$ and retain data $\mathcal{D}_{r}=\mathcal{D}_{train}\setminus\mathcal{D}_{f}$. Let $\mathcal{D}_{test}$ represent test data. $\theta_{u}$ denotes the model after the unlearning process, where the information of forget data $\mathcal{D}_{f}$ has been removed. The purpose of machine unlearning is to enable $\theta_{u}$ to forget $\mathcal{D}_{f}$, while maintaining accuracy on $\mathcal{D}_{r}$ and $\mathcal{D}_{test}$.
@@
Conformal prediction is proposed to quantify uncertainty in machine learning models, providing prediction sets that contain the ground truth label with a theoretically guaranteed probability Angelopoulos and Bates (2021). Among the various types of conformal prediction, this work specifically focuses on split conformal prediction (SCP). To construct the conformal prediction set, it involves four steps on the unlearning model:
@@
    Calibration Data. SCP first chooses unseen data as calibration data, which must be held out from both the training and test sets to ensure independence.
@@
    where $p_{i}({\bm{x}})$ represents the probability of different class $y_{i}$.
@@
    Quantile Computation. Given a target miscoverage rate $\alpha\in[0,1]$, SCP obtains threshold $\hat{q}$ by taking the $1-\alpha$ quantile of the non-conformity score of the ground truth labels $y_{t}$ on the calibration data $({\bm{x}},y_{t})\in\mathcal{D}_{c}$,
    $\hat{q}=\text{Quantile}_{1-\alpha}\big{(}S({\bm{x}},y_{t})\big{)}.$ (2)
@@
2.2 Identifying Fake Unlearning in Existing Unlearning Metrics
Evaluating the effectiveness of machine unlearning involves a trade-off between how well the model forgets data and retains useful knowledge. The first key question we pose is as follows:
(Q1) Can we recover the data that is identified as forgotten by the metrics UA and MIA?
@@ 2.2 Identifying Fake Unlearning in Existing Unlearning Metrics
We propose a conformal prediction-based recovery method. If the ground truth of forget data falls within the conformal prediction set, we consider the recovery successful. Thus, fake unlearning is defined as the scenario where the data identified as forgotten by existing metrics can be recovered by conformal prediction.
To substantiate our claim, we first apply metrics, UA (i.e., $1-$ the accuracy on forget data), retain accuracy (RA, i.e., accuracy on retain data), test accuracy (TA, i.e., accuracy on test data) and MIA, to evaluate 3 classic unlearning methods, Retrain (RT), Finetune (FT) Warnecke et al. (2021), and Random Label (RL) Graves et al. (2021). The results are trained on CIFAR-10 with ResNet-18 in a random data forgetting scenario. In Table 1, the UA and MIA results prove that the models indeed fail to correctly classify some forget data and identify membership. However, higher UA and lower MIA cannot ensure that these forget data points can not appear in any form within the model’s predictions.
@@
We employ conformal prediction to investigate whether we can recover forget data’s ground truth, specifically, whether the ground truth labels still appear within the conformal prediction sets. We set the confidence level to $95\%$, i.e., $\alpha=0.05$, and set the calibration set size to $2000$. In Table 2, we count the number of data points that are identified as truly forgotten by UA and MIA (marked as mis-label) and count how many of these points can still be recovered (marked as in-set). The results of UA reveal that even though the model misclassifies part of the forget data, on average $54.6\%$ of these misclassified data instances are still recovered by conformal prediction. This finding demonstrates that a high UA does not mean the model has truly forgotten the data, and thus relying solely on UA to evaluate the forgetting quality is fragile. A similar phenomenon occurs on results of MIA. In MIA, the prediction of ‘0’ indicates a data point is truly forgotten, while ‘1’ signifies it is still identified as the training membership. The mis-label column of MIA refers to the number of data points that are predicted as ‘0’. The in-set here refers to the number of mis-label data points whose conformal prediction set still includes ‘1’. Thus, the recover ratio indicates that, although the MIA fails to identify an average of $18.33\%$ of the forget data as training membership, conformal prediction can still recover $54.7\%$ of these forget data within prediction sets. For more results of other unlearning methods, see Table 6 in AppendixB.1.
Overall, the high recover ratio observed across both UA and MIA in Tables 2 and 6 highlights that these unlearning methods are not fully effective at eliminating traces of the forget data from the uncertainty quantification perspective and can still be recovered by conformal prediction. This encloses that the fake unlearning issue arises when the ground truth label of misclassified data falls within the conformal prediction set.
@@ 2.3 Designing Metrics Motivated by Conformal Prediction
Based on the limitation of UA and MIA metrics shown in Section 2.2, it raises a question as follows:
(Q2) Can we develop metrics to address the fake unlearning issue of UA and MIA?
@@ 2.3.1 Definition of New Metrics
To overcome the fake unlearning inherent in UA, we introduce a novel metric, CR, which incorporates both coverage and set size in conformal prediction to provide a more comprehensive evaluation. Before defining CR, we introduce Coverage and Set Size.
Given a dataset $\mathcal{D}$, the definition of Coverage is as follows:
@@
where $y_{t}$ is the true label of data point ${\bm{x}}$. Indicator function $\mathbb{I}(\cdot)$ returns $1$ if the enclosed condition is true and 00 otherwise. Coverage reflects the probability that the true label falls within the prediction set ${\mathbb{C}}({\bm{x}})$. For $\mathcal{D}=\mathcal{D}_{f}$, high coverage indicates that the model retains significant information about forget data, suggesting fake unlearning.
Given a dataset $\mathcal{D}$, Set Size is defined as follows:
@@
where $|{\mathbb{C}}({\bm{x}})|$ represents the prediction set size of a specific instance. When $y_{t}\in{\mathbb{C}}({\bm{x}})$, a small set size indicates that fewer non-ground truth classes are included in the prediction set, reflecting stronger fake unlearning.
@@
CR balances the information captured by Coverage and Set Size. A lower CR value implies stronger unlearning on a given dataset. CR is inspired by conformal prediction which is proposed to assess the model’s behavior on new and unseen data, not on the training data. Thus, we emphasize that CR only measures forget data $\mathcal{D}_{f}$ and test data $\mathcal{D}_{test}$.
@@
MIACR is proposed to address the limitation of the existing MIA metric. Among three potential conformal prediction sets {0}, {1} and {0, 1}, only set {0} is ideal case for MIA, because the presence of ‘1’ represents the data point can still be recognised as a training member. Therefore, we introduce a new metric MIACR as:
@@
where ${\mathbb{C}}({\bm{x}})=\{0\}$ denotes prediction set is $\{0\}$. A higher MIACR score indicates a stronger forgetting. While MIA consider a data point forgotten once the logit for label ‘0’ exceeds that for label ‘1’. MIACR applies a stricter rule by also requiring that label ‘1’ be absent from the conformal prediction set, giving a more rigorous check of membership status and forgetting quality.
@@
The existing accuracy-based metrics UA and MIA suffer from a fake unlearning issue, as true labels of misclassified data points may still remain within the prediction set. In contrast, our metrics CR and MIACR eliminate this issue by examining the entire conformal prediction set, yielding a firmer evaluation of forgetting quality and predictive reliability. Besides evidence in Tables 2 and  6, Figures 710 in Appendix also support this claim. The results show the probability of true labels for data recognized as forgotten by CR is consistently closer to 0 when compared to UA, confirming the stricter criterion.
@@
Evaluation Criteria of Our Metrics Similar to existing unlearning metrics, we consider two different criteria [^1] The appropriate evaluation criteria vary across unlearning application scenarios Kurmanji et al. (2023): criterion ➊ is particularly relevant for user privacy scenario, while criterion ➋ focuses on bias removal scenario.to measure unlearning performance with our metrics, Gap to RT Criterion: A lower gap to the RT method is better for both CR and MIACR metrics. The gap relative to RT is represented in blue text (•) in our result tables. Limit-Based Criterion: For the CR, a lower CR value of forget data $\mathcal{D}_{f}$ indicates stronger forgetting performance while a higher CR value of $\mathcal{D}_{test}$ represents higher preserved model utility. For the MIACR, a higher MIACR value for $\mathcal{D}_{f}$ reflects better unlearning effectiveness.
2.3.2 Discussion of Confidence Level and Calibration Set Size